# Open-Vocabulary SAM: Segment and Recognize Twenty-thousand Classes Interactively

### v1

@@ Abstract
The CLIP and Segment Anything Model (SAM) are remarkable vision foundation models (VFMs). SAM excels in segmentation tasks across diverse domains, while CLIP is renowned for its zero-shot recognition capabilities. This paper presents an in-depth exploration of integrating these two models into a unified framework. Specifically, we introduce the Open-Vocabulary SAM, a SAM-inspired model designed for simultaneous interactive segmentation and recognition, leveraging two unique knowledge transfer modules: SAM2CLIP and CLIP2SAM. The former adapts SAM’s knowledge into the CLIP via distillation and learnable transformer adapters, while the latter transfers CLIP knowledge into SAM, enhancing its recognition capabilities. Extensive experiments on various datasets and detectors show the effectiveness of Open-Vocabulary SAM in both segmentation and recognition tasks, significantly outperforming the naïve baselines of simply combining SAM and CLIP. Furthermore, aided with image classification data training, our method can segment and recognize approximately 22,000 classes (examples are shown in Fig. 1).
1 Introduction
@@
While SAM and CLIP offer considerable advantages, they also have inherent limitations in their original designs. SAM, for instance, lacks the capability to recognize the segments it identifies. Efforts to overcome this by integrating a classification head have been made [29, 80], but these solutions are constrained to specific datasets or closed-set settings. On the other hand, CLIP, which is trained using image-level contrastive losses, faces challenges in adapting its representations for dense prediction tasks. To address this, several studies [63, 72, 64, 54, 17] have investigated ways to align CLIP’s representation for dense predictions. However, these approaches tend to be dataset-specific and not universally applicable. For example, some research has focused on open vocabulary segmentation on the ADE-20k [81] dataset, using the COCO [41] dataset for pre-training. Merging SAM and CLIP in a naïve manner, as illustrated in Fig. 2 (a) and (b), proves to be inefficient. This approach not only incurs substantial computational expenses but also yields subpar results, including recognition of small-scale objects, as evidenced by our experimental results.
In this study, we address these challenges with a unified encoder-decoder framework that integrates a CLIP encoder and a SAM decoder, as depicted in Fig. 2 (c). To bridge these two distinct components effectively, we introduce two novel modules, SAM2CLIP and CLIP2SAM, facilitating dual knowledge transfer. First, we distill knowledge from the SAM encoder to a CLIP encoder using SAM2CLIP. This distillation process is uniquely executed not directly on the CLIP encoder, which is kept frozen to maintain its existing knowledge, but rather on a lightweight transformer-like adapter using a pixel-wise distillation loss. The adapter takes multiscale features as input, with the goal of aligning CLIP features with SAM representation. On the decoding side, the CLIP2SAM module transfers knowledge from the frozen CLIP encoder to the SAM decoder. In particular, we design a feature pyramid adapter with a RoIAlign operator to be jointly trained with the SAM decoder.
@@
We conduct extensive experiments across a range of datasets and scenarios, encompassing closed-set and open-vocabulary interactive segmentation. Notably, when compared to basic combined baselines, our approach demonstrates superior performance, achieving over 2% improvement in IoU and 3% in mAP with various detectors on the COCO dataset. In particular, in case of recognition on LVIS, our approach achieves over 20% improvements over previous adapters. Furthermore, by expanding our approach with a more diverse array of datasets, we have developed a versatile, interactive tool suitable for practical applications. For detailed results, we direct the reader to Sec. 4 and the appendix.
@@ 2 Related Work
Open Vocabulary Dense Prediction. This direction aims to recognize region visual concepts of arbitrary categories described by texts, which includes object detection [75, 17, 62, 61, 66], semantic segmentation [68, 83, 34, 84, 37, 35], and panoptic segmentation [67, 72, 71]. This necessitates the alignment between region and text representations with the help of VLMs [23, 50, 55]. For open-vocabulary detection, a series of works [17, 74, 63, 54] distill knowledge from the CLIP models to recognize novel objects. In contrast to distillation-based methods, several works [28, 65] directly build object detectors upon frozen CLIP CNNs. For open-vocabulary segmentation, the typical works [69, 12, 67, 72] first generate class-agnostic mask proposals and then classify the proposals with CLIP. Recently, several works [67, 72] build the mask generator upon the frozen diffusion model [51] and CLIP model. Meanwhile, several studies [49, 48, 45, 24] focus on class-agnostic segmentation and detection to enrich generalization ability in various domains. However, most approaches are trained and tested on specific datasets. Our approach is based on SAM, which provides a general, interactive tool to support different open vocabulary detectors.
Prompting in Computer Vision. Prompting, originating from in-context learning in natural language processing (NLP) as seen in works like Brown et al. [4] and Rubin et al. [52], leverages a large language model to infer unseen tasks through context-specific input-output pairs. Recent studies [3, 58, 59, 85, 15, 1, 43] have explored in-context learning for visual tasks. Common techniques involve mask image modeling [19, 2, 73] for cross-task visual prompting, as employed by approaches like Painter [58] and Bar et al. [3]. SAM [27] demonstrates in-context learning through interactive segmentation, using diverse visual prompts like points, boxes, and masks, although it is limited to class-agnostic mask prediction. Meanwhile, other studies [16, 40, 8, 21, 36] have concentrated on efficient parameter tuning of visual foundation models, typically focusing on a single model. Our work uniquely bridges two models, CLIP and SAM, exploring their combined potential for enhanced general segmentation and recognition capabilities.
Segmentation Anything Model. SAM [27] presents a new data engine and portable model for general object segmentation. Subsequent research has employed SAM as an interactive segmentation tool for various vision tasks, including grounding [42], tracking [10], distillation [77, 82], medical analysis [60, 6], and generation [78]. While most studies use SAM to augment downstream tasks, none have yet integrated VLMs and SAM into a unified model capable of both segmentation and recognition of novel classes. Our work makes the first attempt to merge the capabilities of VLMs with SAM for enhanced task versatility.
@@ 3.1 Preliminaries and Baselines
CLIP. Given an input image and a corresponding caption , the CLIP framework processes these modalities to produce respective embeddings: the image embedding , derived from its image encoder, and the text embedding , obtained from its text encoder. In the context of open-vocabulary object detection and segmentation, CLIP’s capability to generalize beyond fixed class labels is leveraged to replace traditional classifiers. For instance, in open-vocabulary detection scenarios, the text embedding for the -th object category is generated by inputting the category name into the CLIP text encoder. This process can employ a single template prompt, such as ”a photo of {category},” or multiple prompt templates. Subsequently, for a given region embedding , that is produced by the RoI-Align [20], the classification score for the -th category is computed as follows:
where denotes the cosine similarity, and is a learnable or fixed temperature to re-scale the value.
Combined Baselines. We introduce two different baselines for combining CLIP and SAM, as depicted in Fig. 2 (a) and (b). The first approach, termed the ‘cropped image baseline’, employs the SAM mask decoder’s output to segment and resize the original input image. This processed image then serves as the input for the CLIP image encoder, and, in conjunction with the CLIP text embedding, the mask is classified using Equ. (1). The second approach, referred to as the ‘cropped CLIP image feature baseline’, employs the same initial CLIP feature extraction step. However, in this method, masks predicted by the SAM decoder are used to crop the CLIP image features. Subsequent pooling of these masked features yields the final label, akin to baseline (a).
@@ 3.2 Open Vocabulary SAM
SAM2CLIP. To resolve , we design the SAM2CLIP module that bridges the gap in feature representations learned by SAM and CLIP, using adaptation and distillation methods. Through comprehensive experiments, we discovered that employing distillation loss along with transformer-based adapters [13], yields effective results. Specifically, the distillation process involves a simple pixel-wise approach, where SAM-Huge serves as the teacher and the frozen CLIP equipped with an adapter assumes the student’s role. We then implement a per-pixel mean squared error (MSE) loss to align the SAM feature with the CLIP feature , as detailed below:
@@
where comprises several transformer layers, and is achieved by bilinear upsampling and addition. With SAM2CLIP, we can even achieve comparable segmentation results with the SAM-Huge on with much lower computational costs. As detailed in Sec. 4.2, we observe that employing convolution-based methods specifically designed for backbone adaptation [14, 8] results in sub-optimal outcomes. The reason for this might be in the inherent architecture of the SAM encoder, which is purely based on ViT. A symmetrical structure is crucial for effective knowledge transfer. With the implementation of SAM2CLIP, we are able to achieve segmentation results comparable to those of SAM-Huge across various detectors, while significantly reducing computational costs.
| Method | COCO | LVIS | FLOPs (G) | #Params (M) | ||||
@@
| Image-Crop baseline | 78.1 | 81.4 | 46.2 | 78.3 | 81.6 | 9.6 | 3,748 | 808 |
| Feature-Crop baseline | 78.1 | 81.4 | 55.1 | 78.3 | 81.6 | 26.5 | 3,545 | 808 |
| Image-Crop baseline + Adapter [85] | 79.6 | 82.1 | 62.0 | 80.1 | 82.0 | 32.1 | 3,748 | 808 |
| Feature-Crop baseline + Adapter [85] | 79.6 | 82.1 | 70.9 | 80.1 | 82.0 | 48.2 | 3,545 | 808 |
| Feature-Crop baseline + Adapter [8] | 78.5 | 80.9 | 75.3 | 79.5 | 81.8 | 29.6 | 3,590 | 828 |
| Feature-Crop baseline + Adapter [14] | 76.1 | 79.3 | 63.1 | 76.1 | 79.2 | 21.0 | 3,580 | 812 |
| Open-Vocabulary SAM | 81.5 | 84.0 | 84.3 | 80.4 | 83.1 | 66.6 | 1,180 | 304 |
| Feature-Crop baseline* | 60.7 | 66.7 | 32.1 | 53.0 | 62.3 | 11.0 | 3,545 | 808 |
| Feature-Crop baseline + Adapter [85]* | 64.7 | 66.7 | 35.1 | 58.9 | 64.2 | 13.2 | 3,545 | 808 |
| Open-Vocabulary SAM* | 68.4 | 65.2 | 76.7 | 63.6 | 67.9 | 60.4 | 1,180 | 304 |
CLIP2SAM. This module aims to leverage CLIP’s knowledge to enhance the recognition capabilities of the SAM decoder. A straightforward approach involves appending a label token to the existing mask token and IoU token . Using , we introduce a specialized adapter to facilitate the transfer of knowledge from the frozen CLIP to the SAM decoder. Subsequently, the enhanced , combined with the output of the prompt encoder and adapted CLIP features, is fed into a two-way transformer. Following the cross-attention process, the improved undergoes further refinement through a multilayer perceptron (MLP), ensuring better alignment with CLIP’s text embedding. The final labels are derived by calculating the distance between the refined label token and the CLIP text embedding, as in Equ. (1).
Open Vocabulary. To tackle , the open-vocabulary challenge, we leverage the knowledge embedded in the frozen CLIP backbone, which aids in recognizing novel and unseen objects during inference. In line with previous studies [28, 64], we fuse the learned class scores with those from the frozen CLIP via a geometric mean to leverage information from both the CLIP backbone and CLIP2SAM. Additionally, we investigate various strategies to expand the vocabulary size, such as joint training with multiple datasets, as detailed in Sec. 4.2. Our experimental results demonstrate that the model scales effectively with large datasets.
3.3 Training and Application
Training and Loss Function. We first use the SAM-1B (1%) dataset [27] for training the SAM2CLIP module to transfer SAM’s knowledge into open-vocabulary SAM, with the loss (Equ. (3)). Then, we joint train the CLIP2SAM and mask decoder using segmentation mask and label annotations from COCO or LVIS. The final loss function is given as . Here, is the Cross-Entropy (CE) loss for mask classification, and and are mask Cross Entropy (CE) loss and Dice loss [46] for segmentation, respectively. In addition, we adopt joint training with the ImageNet dataset for our Open-Vocabulary SAM for demo (See Fig. 5).
Inference and Demo Tools. Our model performs inference like SAM, with points and boxes as visual prompts. Specifically, we test boxes and points as visual prompts for the encoder in Sec. 4. On the project page, we show a demo of our model, which can segment and recognize with prompts.
4 Experiment
Datasets and Metrics. We mainly use COCO [41] and LVIS [18] datasets for the experiments. Moreover, we also use part of SAM data (1%) for SAM2CLIP knowledge transfer. For COCO, we report the results of both close-set and open-vocabulary settings for the instance segmentation task. In particular, following Zareian et al. [75], we split 48 base classes with annotations and 17 target classes without annotations. We use the base class annotations for training. For LVIS datasets, we adopt the open-vocabulary setting and report the results of for novel classes. In addition, we also report the accuracy of each box or point prompt for reference since our goal is to add recognition ability to SAM. Meanwhile, each prompt’s intersection-over-union (IoU) with its ground truth mask is also adopted to verify the segmentation ability of our method. Different from previous open vocabulary segmentation tasks, where the proposals are generated by the detectors themselves, we term our setting as open-vocabulary interactive segmentation, where the boxes or points prompts serve as conditional inputs.
@@ 4 Experiment
Implementation Details. We implement our models in PyTorch [47] with both MMDetection framework [7] and SAM codebase [27]. We use 8 A100 GPUs for distributed training. Each mini-batch has two images per GPU. The optimizer is AdamW [44] with a weight decay of 0.0001. We adopt full image size for a random crop in the pre-training and training process following Cheng et al. [9]. All the class names are transferred into CLIP text embedding, following previous works [17]. We train each model for 12 epochs for fair comparison. Due to the limitation of computation costs, we do not adopt joint SAM data and COCO data training. We first perform training the SAM2CLIP on SAM data, and then we finetune the model on COCO or LVIS data. Please refer to the supplementary material for more details.
4.1 Main Results
| Method | Detectors | mAP | AP50 | AP75 | APS | APM | APL | #Params (M) | FLOPs (G) |
| SAM-Huge | Faster-RCNN (R50) | 35.6 | 54.9 | 38.4 | 17.2 | 39.1 | 51.4 | 641 | 3,001 |
| SAM-Huge (finetuned) | Faster-RCNN (R50) | 35.8 | 55.0 | 38.4 | 16.5 | 38.6 | 53.0 | 641 | 3,001 |
| Open-Vocabulary SAM | Faster-RCNN (R50) | 35.8 | 55.6 | 38.3 | 16.0 | 38.9 | 53.1 | 304 | 1,180 |
| SAM-Huge | Detic (swin-base) | 36.4 | 57.1 | 39.4 | 21.4 | 40.8 | 54.6 | 641 | 3,001 |
| SAM-Huge (finetuned) | Detic (swin-base) | 36.8 | 57.4 | 39.8 | 20.8 | 40.6 | 55.1 | 641 | 3,001 |
| Open-Vocabulary SAM | Detic (swin-base) | 36.7 | 57.2 | 39.7 | 20.7 | 40.8 | 54.9 | 304 | 1,180 |
Comparison with Combined Baselines Using Ground Truth. To avoid the influence of other modules, we first demonstrate the recognition ability of our model in Tab. 1. Compared to the simple combined approaches, adding various adapters with joint co-training leads to better results. However, the recognition ability is still limited on both COCO and LVIS. Our Open-Vocabulary SAM achieves the best results on both boxes and points as visual prompts. We observe more significant gains on LVIS datasets. We argue that LVIS contains more small objects, which is more challenging than COCO. Our method can solve and lead to over 20% accuracy improvement. Although the segmentation quality is pretty good (about 80 IoU on COCO and LVIS with box prompt), our method still achieves 2% IoU improvements. This indicates the effectiveness of our joint co-training on mask prediction and classification. Compared with boxes as prompts, using points as prompts is more challenging since the location clues of points are much weaker than boxes. However, our approach is still better than combined baselines or them with adapters.
Comparison with Combined Baselines on OV-Detector. In Tab. 2, we adopt a more challenging setting by using the box prediction from the existing open-vocabulary detector to simulate the interactive segmentation process with deviation. We choose the representative Detic [86] as the open-vocabulary detector. Again, our method also achieves the best performance on both COCO and LVIS datasets. In particular, on COCO, compared with previous works [85], our method achieves 3.0 mask mAP improvements with much lower parameter costs. Results with more detectors can be found in the supplementary.
Comparison with SAM on various detectors. In Tab. 3, we also test the mask prediction quality of our model and original SAM on two different detectors. Our method can achieve better performance than the original SAM and perform comparably with SAM fine-tuned on COCO. It is worth noting that our Open-Vocabulary SAM has much lower computational costs and parameters than SAM.
Visualization Comparison. In Fig. 4, we compare our approach with the feature-crop baseline. Our model shows a better performance in classifying small and rare object classification, as well as handling occlusion scenarios.
Model as a Zero Shot Annotation Tool. In addition to COCO and LVIS standard datasets training, following the spirit of SAM, we also scale up our model by training it with more data. In particular, we adopt more detection data (V3Det [57], Object365 [53]) and classification data (ImageNet22k [11]). Owing to significant costs, we have not conducted comparisons with other baselines for this setting. Rather, we have adapted our method into an interactive annotation tool capable of segmenting and recognizing over 22,000 classes.
4.2 Ablation Studies and Analysis

### v2

# Project Page: https://www.mmlab-ntu.com/project/ovsam

@@ Abstract
The CLIP and Segment Anything Model (SAM) are remarkable vision foundation models (VFMs). SAM excels in segmentation tasks across diverse domains, whereas CLIP is renowned for its zero-shot recognition capabilities. This paper presents an in-depth exploration of integrating these two models into a unified framework. Specifically, we introduce the Open-Vocabulary SAM, a SAM-inspired model designed for simultaneous interactive segmentation and recognition, leveraging two unique knowledge transfer modules: SAM2CLIP and CLIP2SAM. The former adapts SAM’s knowledge into the CLIP via distillation and learnable transformer adapters, while the latter transfers CLIP knowledge into SAM, enhancing its recognition capabilities. Extensive experiments on various datasets and detectors show the effectiveness of Open-Vocabulary SAM in both segmentation and recognition tasks, significantly outperforming the naïve baselines of simply combining SAM and CLIP. Furthermore, aided with image classification data training, our method can segment and recognize approximately 22,000 classes.
Keywords:
Scene Understanding Promptable Segmentation1 Introduction
@@
While SAM and CLIP offer considerable advantages, they also have inherent limitations in their original designs. SAM, for instance, lacks the capability to recognize the segments it identifies. Efforts to overcome this by integrating a classification head have been made [32, 84], but these solutions are constrained to specific datasets or closed-set settings. On the other hand, CLIP, which is trained using image-level contrastive losses, faces challenges in adapting its representations for dense prediction tasks. To address this, several studies [66, 76, 67, 57, 18] have investigated ways to align CLIP’s representation for dense predictions. However, these approaches tend to be dataset-specific and not universally applicable. For example, some research has focused on open vocabulary segmentation on the ADE-20k [85] dataset, using the COCO [43] dataset for pre-training. Merging SAM and CLIP in a naïve manner, as illustrated in Fig. 2 (a) and (b), proves to be inefficient. This approach incurs substantial computational expenses and yields subpar results, including recognition of small-scale objects, as evidenced by our experimental results.
In this study, we address these challenges with a unified encoder-decoder framework that integrates a CLIP encoder and a SAM decoder, as depicted in Fig. 2 (c). To bridge these two distinct components effectively, we introduce two novel modules, SAM2CLIP and CLIP2SAM, facilitating dual knowledge transfer. First, we distill knowledge from the SAM encoder to the CLIP encoder using SAM2CLIP. This distillation process is uniquely executed not directly on the CLIP encoder, which is kept frozen to maintain its existing knowledge, but rather on a lightweight transformer-like adapter using a pixel-wise distillation loss. The adapter takes multi-scale features as input, with the goal of aligning CLIP features with SAM representation. On the decoding side, the CLIP2SAM module transfers knowledge from the frozen CLIP encoder to the SAM decoder. In particular, we design a feature pyramid adapter with a RoIAlign operator to be jointly trained with the SAM decoder. Both modules are lightweight and naturally combine the strengths of CLIP and SAM.
@@
We conduct extensive experiments across a range of datasets and scenarios, encompassing closed-set and open-vocabulary interactive segmentation. Notably, when compared to basic combined baselines, our approach demonstrates superior performance, achieving over 2% improvement in IoU and 3% in mAP with various detectors on the COCO dataset. In particular, in the case of recognition on LVIS, our approach achieves over 20% improvements over previous adapters. Furthermore, by expanding our approach with a more diverse array of datasets, we have developed a versatile, interactive tool suitable for practical applications. For detailed results, we direct the reader to Sec. 4 and the appendix.
@@ 2 Related Work
Open Vocabulary Dense Prediction. This direction aims to recognize region visual concepts of arbitrary categories described by texts, which includes object detection [79, 18, 64, 63, 69], semantic segmentation [71, 87, 37, 88, 40, 38], and panoptic segmentation [70, 76, 75]. This necessitates the alignment between region and text representations with the help of VLMs [26, 53, 58]. For open-vocabulary detection, a series of works [18, 78, 66, 57] distill knowledge from the CLIP models to recognize novel objects. In contrast to distillation-based methods, several works [31, 68] directly build object detectors upon frozen CLIP CNNs. For open-vocabulary segmentation, the typical works [72, 12, 70, 76] first generate class-agnostic mask proposals and then classify the proposals with CLIP. Recently, several works [70, 76] build the mask generator upon the frozen diffusion model [54] and CLIP model [53]. Meanwhile, several studies [52, 51, 47, 27] focus on class-agnostic segmentation and detection to enrich generalization ability in various domains. However, most approaches are trained and tested on specific datasets. Our approach is based on SAM, which provides a general, interactive tool to support different open vocabulary detectors.
Prompting in Computer Vision. Prompting, originating from in-context learning in natural language processing (NLP) as seen in works like Brown et al. [4] and Rubin et al. [55], leverages a large language model to infer unseen tasks through context-specific input-output pairs. Recent studies [3, 61, 62, 89, 16, 1, 45] have explored in-context learning for visual tasks. Common techniques involve mask image modeling [21, 2, 77] for cross-task visual prompting, as employed by approaches like Painter [61] and Bar et al. [3]. SAM [30] demonstrates in-context learning through interactive segmentation, using diverse visual prompts like points, boxes, and masks, although it is limited to class-agnostic mask prediction. Meanwhile, other studies [17, 42, 8, 23, 39] have concentrated on efficient parameter tuning of visual foundation models, typically focusing on a single model. Our work uniquely bridges two models, CLIP and SAM, exploring their combined potential for enhanced general segmentation and recognition capabilities. In particular, we adopt visual prompts (box, point) as the model’s inputs.
Segmentation Anything Model. SAM [30] presents a new data engine and portable model for segmentation. Subsequent research has employed SAM as an interactive segmentation tool for various vision tasks, including grounding [44], tracking [10], distillation [81, 86], medical analysis [65, 5], and generation [82]. While some studies use SAM and CLIP for segmentation [6, 20, 59, 36, 49], none have yet integrated VLMs and SAM into a unified model capable of both segmentation and recognition of novel classes. Our work makes the first attempt to merge the capabilities of VLMs with SAM for enhanced task versatility.
@@ 3.1 Preliminaries and Baselines
CLIP. Given an input image and a corresponding caption , the CLIP framework processes these modalities to produce respective embeddings: the image embedding , derived from its image encoder, and the text embedding , obtained from its text encoder. In the context of open-vocabulary object detection and segmentation, CLIP’s capability to generalize beyond fixed class labels is leveraged to replace traditional classifiers. For instance, in open-vocabulary detection scenarios, the text embedding for the -th object category is generated by inputting the category name into the CLIP text encoder. This process can employ a single template prompt, such as "a photo of {category}," or multiple prompt templates. Subsequently, for a given region embedding , that is produced by the RoI-Align [22], the classification score for the -th category is computed as: , where denotes the cosine similarity, and is a learnable or fixed temperature to re-scale the value.
Combined Baselines. We introduce two different baselines for combining CLIP and SAM, as depicted in Fig. 2 (a) and (b). The first approach, termed the ‘cropped image baseline’, employs the SAM mask decoder’s output to segment and resize the original input image. This processed image then serves as the input for the CLIP image encoder, and, in conjunction with the CLIP text embedding, the mask is classified using the similarities between visual and text embeddings. The second approach referred to as the ‘cropped CLIP image feature baseline’, employs the same initial CLIP feature extraction step. However, in this method, masks predicted by the SAM decoder are used to crop the CLIP image features. Subsequent pooling of these masked features yields the final label, akin to baseline (a).
@@ 3.2 Open Vocabulary SAM
SAM2CLIP. To resolve , we design the SAM2CLIP module that bridges the gap in feature representations learned by SAM and CLIP, using adaptation and distillation methods. Through comprehensive experiments, we discovered that employing distillation loss along with transformer-based adapters [14], yields effective results. Specifically, the distillation process involves a simple pixel-wise approach, where SAM-Huge serves as the teacher, and the frozen CLIP equipped with an adapter assumes the student’s role. We then implement a per-pixel mean squared error (MSE) loss to align the SAM feature with the CLIP feature , as detailed below:
@@
where comprises several transformer layers, and is achieved by bilinear upsampling and addition.
With SAM2CLIP, we can even achieve comparable segmentation results with the SAM-Huge with much lower computational costs. As detailed in Sec. 4.2, we observe that employing convolution-based methods specifically designed for backbone adaptation [15, 8] results in sub-optimal outcomes. The reason for this might be in the inherent architecture of the SAM encoder, which is purely based on ViT. A symmetrical structure is crucial for effective knowledge transfer. With the implementation of SAM2CLIP, we can achieve segmentation results comparable to those of SAM-Huge across various detectors, while significantly reducing computational costs.
CLIP2SAM. This module aims to leverage CLIP’s knowledge to enhance the recognition capabilities of the SAM decoder. A straightforward approach involves appending a label token to the existing mask token and IoU token . Using , we introduce a specialized adapter to facilitate the transfer of knowledge from the frozen CLIP to the SAM decoder. Subsequently, the enhanced , combined with the output of the prompt encoder and adapted CLIP features, is fed into a two-way transformer. Following the cross-attention process, the improved undergoes further refinement through a multilayer perceptron (MLP), ensuring better alignment with CLIP’s text embedding. The final labels are derived by calculating the distance between the refined label token and the CLIP text embedding.
Open Vocabulary. To tackle , the open-vocabulary challenge, we leverage the knowledge embedded in the frozen CLIP backbone, which aids in recognizing novel and unseen objects during inference. In line with previous studies [31, 67], we fuse the learned class scores with those from the frozen CLIP via a geometric mean to leverage information from both the CLIP and CLIP2SAM. Additionally, we investigate various strategies to expand the vocabulary size, such as joint training with multiple datasets, as detailed in Sec. 4.2. Our experimental results show that the model scales effectively with large datasets.
While our approach can address the open vocabulary challenge, it is important to distinguish the setting of Open-Vocabulary SAM from that of previous open vocabulary segmentation methods [6, 24, 13, 74]. Unlike previous techniques, which typically depend on a separate segmentor to produce mask proposals, our method only uses visual prompts, such as boxes and points, to generate masks and labels. Furthermore, our method specifically targets foreground objects, aligning with the ImageNet [11] datasets and the purpose of box prompts. To show the effectiveness of our method, we compare the performance of our method with previous open vocabulary segmentation methods in Sec. 4.1.
3.3 Training and Application
Training and Loss Function. We first use the SA-1B (1%) dataset [30] for training the SAM2CLIP module to transfer SAM’s knowledge into open-vocabulary SAM, with the loss (Equ. (2)). Then, we joint train the CLIP2SAM and mask decoder using segmentation mask and label annotations from COCO or LVIS. The final loss function is given as . Here, is the Cross-Entropy (CE) loss for mask classification, and and are mask Cross Entropy (CE) loss and Dice loss [48] for segmentation, respectively. In addition, we adopt joint training with the ImageNet dataset for classifying over 22,000 categories.
Inference and Demo Tools. Our model performs inference like SAM, with points and boxes as visual prompts. Specifically, we test boxes and points as visual prompts for the encoder in Sec. 4. In the appendix, we show a demo of our model, which can segment and recognize with prompts.
4 Experiments
Datasets and Metrics. We mainly use COCO [43] and LVIS [19] datasets for the experiments. Moreover, we also use part of SA-1B [30] data (1%) for SAM2CLIP knowledge transfer. For COCO, we report the results of both close-set and open-vocabulary settings for the instance segmentation task. In particular, following Zareian et al. [79], we split 48 base classes with annotations and 17 target classes without annotations. We use the base class annotations for training. For LVIS datasets, we adopt the open-vocabulary setting and report the results of for novel classes. For evaluation metrics, we report the accuracy of recognition for reference to evaluate the recognition capability. Meanwhile, each prompt’s intersection-over-union (IoU) with its ground truth mask is also adopted to evaluate the segmentation ability of our method. As mentioned in Sec. 3.2, different from previous open vocabulary segmentation tasks, boxes or points serve as visual prompts in our method.
Method | COCO | LVIS | FLOPs | #Param | ||||
@@ 4 Experiments
Image-Crop baseline | 78.1 | 81.4 | 46.2 | 78.3 | 81.6 | 9.6 | 3,748G | 808M |
Feature-Crop baseline | 78.1 | 81.4 | 55.1 | 78.3 | 81.6 | 26.5 | 3,545G | 808M |
Image-Crop baseline + CoOp [89] | 79.6 | 82.1 | 62.0 | 80.1 | 82.0 | 32.1 | 3,748G | 808M |
Feature-Crop baseline + CoOp [89] | 79.6 | 82.1 | 70.9 | 80.1 | 82.0 | 48.2 | 3,545G | 808M |
Open-Vocabulary SAM | 81.5 | 84.0 | 84.3 | 80.4 | 83.1 | 66.6 | 1,180G | 304M |
Image-Crop baseline* | 60.7 | 66.7 | 24.5 | 53.0 | 62.3 | 6.2 | 3,748G | 808M |
Feature-Crop baseline* | 60.7 | 66.7 | 32.1 | 53.0 | 62.3 | 11.0 | 3,545G | 808M |
Image-Crop baseline + CoOp [89]* | 64.7 | 66.7 | 28.2 | 58.9 | 64.2 | 8.3 | 3,748G | 808M |
Feature-Crop baseline + CoOp [89]* | 64.7 | 66.7 | 35.1 | 58.9 | 64.2 | 13.2 | 3,545G | 808M |
Open-Vocabulary SAM* | 68.4 | 65.2 | 76.7 | 63.6 | 67.9 | 60.4 | 1,180G | 304M |
@@
Implementation Details. We implement our models in PyTorch [50] with MMDetection [7]. We use 8 A100 GPUs for distributed training. Each mini-batch has two images per GPU. The optimizer is AdamW [46] with a weight decay of 0.0001. We adopt full image size for a random crop in the pre-training and training process following Cheng et al. [9]. All the class names are transferred into CLIP text embedding, following previous works [18]. We train each model for 12 epochs for fair comparison. Due to the limitation of computation costs, we do not adopt joint SAM data and COCO data training. We first perform training the SAM2CLIP on SA-1B (1%), and then we finetune the model on COCO or LVIS data. Please refer to the supplementary material for more details.
Benchmark Setup. Open Vocabulary SAM is different from both SAM [30] and open vocabulary segmentation methods [24, 13, 74]. Compared with Open-Vocabulary SAM, SAM [30] cannot recognize segmented objects while open vocabulary segmentation methods [24, 13, 74] usually rely on a separate segmentor (e.g., Mask-RCNN [22]) instead of prompts like points or boxes. For a complete evaluation of Open-Vocabulary SAM, we set the benchmark as three parts, including 1).comparison with combined baselines to verify the effectiveness; 2).comparison with open-vocabulary segmentation methods to evaluate the recognition; 3).comparision with SAM models to evaluate segmentation.
Method | Detectors | mAP | AP50 | AP75 | APS | APM | APL | #Params | FLOPs |
SAM-Huge | Faster-RCNN (R50) | 35.6 | 54.9 | 38.4 | 17.2 | 39.1 | 51.4 | 641M | 3,001G |
SAM-Huge (finetuned) | Faster-RCNN (R50) | 35.8 | 55.0 | 38.4 | 16.5 | 38.6 | 53.0 | 641M | 3,001G |
Open-Vocabulary SAM | Faster-RCNN (R50) | 35.8 | 55.6 | 38.3 | 16.0 | 38.9 | 53.1 | 304M | 1,180G |
SAM-Huge | Detic (swin-base) | 36.4 | 57.1 | 39.4 | 21.4 | 40.8 | 54.6 | 641M | 3,001G |
SAM-Huge (finetuned) | Detic (swin-base) | 36.8 | 57.4 | 39.8 | 20.8 | 40.6 | 55.1 | 641M | 3,001G |
Open-Vocabulary SAM | Detic (swin-base) | 36.7 | 57.2 | 39.7 | 20.7 | 40.8 | 54.9 | 304M | 1,180G |
SAM-Huge | ViTDet (Huge) | 46.3 | 72.0 | 49.8 | 25.2 | 45.5 | 59.6 | 641M | 3,001G |
SAM-Huge (finetuned) | ViTDet (Huge) | 46.5 | 72.3 | 50.3 | 25.2 | 45.8 | 60.1 | 641M | 3,001G |
Open-Vocabulary SAM | ViTDet (Huge) | 48.8 | 73.8 | 52.9 | 24.8 | 46.3 | 64.2 | 304M | 1,180G |
Datasets | Accuracy | #vocaulary | #images |
LVIS | 83.1 | 1,203 | 99K |
V3Det | 78.7 | 13,204 | 183K |
I-21k | 44.5 | 19,167 | 13M |
V3Det + LVIS | 82.7 | 13,844 | 282K |
V3Det + LVIS + I-21k | 83.3 | 25,898 | 13M |
V3Det + LVIS + I-21k + Object365 | 83.0 | 25,970 | 15M |